# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix G Experiment Results of Detecting 2024 Olympic News or Machine-generated News
Results of Tests in Two Cases. Figure 5 and 6 show the results of detecting real Olympic news or news generated by Gemini-1.5-Flash , Gemini-1.5-Pro and PaLM 2 with designating human-written text from XSum dataset as $x_{t}$ in Scenario 1 and Scenario 2 respectively. In Scenario 1, our algorithm consistently controls the FPRs below the significance level $\alpha$ for all source models, scoring models and score functions. It is because we used the real $\Delta$ value between two sequences of human texts as $\epsilon$, which satisfies the condition of $H_{0}$, i.e., $\lvert\mu_{x}-\mu_{y}\rvert\leq\epsilon$. This ensures that the wealth remains a supermartingale. Texts generated by PaLM 2 are detected almost immediately by most score functions within 100 time steps as illustrated in Figure 5(e) and Figure 5(f). Conversely, fake Olympic news generated by Gemini-1.5-Pro often fails to be identified as LLM-generated before 500 by some score functions, as shown in Figure 5(c) and Figure 5(d). Vertical lines in Figure 5(a)-5(d) are displayed because the the $\Delta$ values for human texts and fake news, as shown in Table 1, are smaller than the corresponding $\epsilon$ values. This means that the score discrepancies between fake news and XSum texts do not exceed the threshold necessary for rejecting $H_{0}$. Although under $H_{1}$, the $\Delta$ for Entropy when using Gemma-2B to score texts generated by Gemini-1.5-Pro is 0.2745, larger than the value of $\epsilon$ ($0.2690$), the discrepancy is too small to lead to a rejection of $H_{0}$ within 500 time steps.
@@
(a) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(b) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(c) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(d) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(e) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
Figure 5: Results for detecting 2024 Olympic news and machine-generated news with our algorithm for Scenario 1. We use 3 source models: Gemini-1.5-Flash, Gemini-1.5-Pro and PaLM 2 to generate fake news and 2 scoring models: Neo-2.7, Gemma-2B. The left column displays results using the Neo-2.7 scoring model, while the right column presents results using the Gemma-2B scoring model. Score functions of supervised classifiers (RoB-base and RoB-large) are independent of scoring models.
(a) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(b) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(c) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(d) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(e) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
Table 1: Values of $\Delta$, which are calculated according to $\Delta=\left\lvert\left({\sum_{i=1}^{500}\phi(x_{i})}\right)/{500}-\left({\sum_{j=1}^{500}\phi(y_{j})}\right)/{500}\right\rvert$, where $x_{i}$ is score of the $i$-th text from XSum, $y_{j}$ is score of the $j$-th text from the source to be detected. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ for each test scenario. For instance, in calculating $\Delta$ for the third column, $y_{j}$ represents the $j$-th fake news generated by Gemini-1.5-Flash based on pre-tokens of Olympic 2024 news. For the fourth column, $y_{j}$ refers to the $j$-th 2024 Olympic news articles. Values in Column “Human, Human" are also used to set $\epsilon$ values for tests in Scenario 1.
@@
Table 2: Values of $d_{t}$ used in Scenario 1, for which we assume that the range of $g_{t}=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ is known beforehand. Specifically, $d_{t}$ is calculated as $\max_{i,j\leq 500}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by Gemini-1.5-Flash, prompted by pre-tokens of 2024 Olympic news. This calculation ensures that $d_{t}\geq\lvert\phi(x_{t})-\phi(y_{t})\rvert$ for any time point $1\leq t\leq 500$. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column Similarly, the fourth column calculates the maximum difference between the scores of all XSum texts and texts sample of 2024 Olympic news.. The derived $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},1/2d_{t}]$.
@@
Moreover, we found that the rejection time is related to the relative magnitude of $(\Delta-\epsilon)$ and $(d_{t}-\epsilon)$. According to the definition of nonnegative wealth $W_{t}^{A}=W_{t-1}^{A}(1-\theta_{t}(g_{t}-\epsilon))$ or $W_{t}^{B}=W_{t-1}^{B}(1-\theta_{t}(-g_{t}-\epsilon))$, large $(-\theta_{t})$ within the range $[0,1/2d_{t}]$ will result in large wealth which allows to quickly reach the threshold for wealth to correctly declare the unknown source as an LLM. Based on the previous proposition of the expected time upper bound for composite hypothesis, we guess that the actual rejection time in our experiment is probably related to the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$, where $d_{t}$ is a certain value for any $t$ in each test as shown in Table 2. We define the relative magnitude as $(\Delta-\epsilon)/(d_{t}-\epsilon)$ and sort the score functions by this ratio from largest to smallest for Scenario 1, as displayed in the Rank column in Table 3. This ranking roughly corresponds to the chronological order of rejection shown in in Figure 5. The quick declaration of an LLM source when $y_{t}$ is generated by PaLM 2 and the slower rejection of $H_{0}$ when $y_{t}$ is generated by Gemini-1.5-Pro in Figure 5 can thus be attributed to the relatively larger and smaller values of $(\Delta-\epsilon)/(d_{t}-\epsilon)$, respectively. The negative ratios attributes to the reason that $\Delta<\epsilon$, which result in the vertical lines in figures. Similarly, we also show the values of $(\Delta-\epsilon)/(d_{t}-\epsilon)$ for Scenario 2 as shown in Table 4, the conclusions in Scenario 1 still holds true for Scenario 2.
@@
Table 3: Values of the ratio $(\Delta-\epsilon)/(d_{t}-\epsilon)$ for Scenario 1, where $\Delta$ and $\epsilon$ are listed in Table 1, $d_{t}$ are shown in Table 2. We sort the score function according to the ratio from largest to smallest, as shown in the Rank column. This ranking roughly corresponds to the chronological order of rejection in in Figure 5.
@@
Table 4: Values of the ratio $(\Delta-\epsilon)/(d_{t}-\epsilon)$ for Scenario 2, where $\Delta$ and $\epsilon$ are listed in Table 1 and Table 5 respectively, $d_{t}$ are shown in Table 6. We sort the scoring function according to the ratio from largest to smallest, as shown in the Rank column. This ranking roughly corresponds to the chronological order of rejection in in Figure 6.
@@
Table 5: Average values of $\epsilon$ used in Scenario 2 estimated by 20 texts in sequence of human-written text $x_{t}$. Every two columns starting from the third column represent the $\epsilon$ values used for $H_{1}$ and $H_{0}$ for each test scenario. For instance, the third column calculates $\epsilon$ for tests between XSum text and Gemini-1.5-Flash-generated text sequences by scoring 20 XSum texts, dividing them into two equal groups, and then doubling the average absolute mean difference between these groups across 1000 random shuffles. The fourth column follows the same method to determine the $\epsilon$ value for tests between XSum texts and 2024 Olympic news. This is calculated as $\epsilon=2\cdot\frac{1}{1000}\sum_{n=1}^{1000}\left\lvert\left({\sum_{i=1}^{10}\phi(x_{i}^{(n)}})\right)/{10}-\left({\sum_{i=11}^{20}\phi(x_{i}^{(n)}})\right)/{10}\right\rvert$, where $\phi(x_{i}^{(n)})$ denotes the score of the $i$-th text after the $n$-th random shuffling of 20 text scores.
@@
Table 6: Average values of $d_{t}$ used in Scenario 2 estimated by previous 10 texts of each sequence. Every two columns starting from the third column represent the $d_{t}$ values used for $H_{1}$ and $H_{0}$ for each test scenario. Specifically, for the third column, we get first 10 samples $x_{i\geq 10}$ from XSum and first 10 observed texts $y_{j\geq 10}$ generated by Gemini-1.5-Flash. We then calculate the maximum difference between $\phi(x_{i})$ and $\phi(y_{j})$ for any $1\leq i\leq 10$, $1\leq j\leq 10$. We double this maximum value to estimate $d_{t}$ value, i.e., $d_{t}=2\cdot\max_{i,j\leq 10}\lvert\phi(x_{i})-\phi(y_{j})\rvert$. The fourth column follows a similar calculation for detecting 2024 Olympic news with XSum texts, where $\phi(x_{i})$ represents score of $i$-th text from XSum, $\phi(y_{j})$ denotes the score of the $j$-th text of 2024 Olympic news.
@@
Based on the results, when we know the actual value of $d_{t}$ and $\epsilon$, our algorithm is very effective. When we estimate their values based on the previous samples that we get, the algorithm can still exhibit a good performance for most score functions. It can be inferred that the rejection time and FPRs of Algorithm 1 are effected by the score function $\phi(\cdot)$ and the scoring model that we select for our algorithm. If the configuration can further amplify the score discrepancy between human-written texts and machine-generated texts, the rejection time can be shortened. If the the score discrepancy between human texts is small, the value FPR will be low.
Comparisons with the Baselines. Permutation test is a fixed-time test. Our goal is to test whether the source of text $y_{t}$ is the same as that of the human-written text $x_{t}$, i.e., whether their scores are from the same distribution. If we choose the mean value as the test statistic, the null hypothesis is that the means are equal ($H_{0}:\mu_{x}=\mu_{y}$) with a batch size of k. Once we have generated $k$ samples from these two sources, we conduct the test. Under the assumption that $H_{0}$ is true, the samples are drawn from the same distribution, which means that the observed discrepancy between the two batches of scores is supposed to be minimal. The test determines whether this difference between the sample means is large enough to reject at a significance level. Specifically, we compare the p-value with the significance level: $\alpha$ for each batch in an uncorrected test, and $\alpha/2^{j}$ for the $j$-th batch in a corrected test. The permutation test is conducted as below:
(1) Calculate the observed mean difference $\Delta=\lvert\sum_{i=1}^{k}\phi(x_{i})/k-\sum_{i=1}^{k}\phi(y_{i})/k\rvert$ of these two batches, and assume that $H_{0}$ is true;
(2) Combine these two sequences into one dataset, reshuffle the data, and divide it into two new groups. This is the permutation operation. Calculate the sampled absolute mean difference for the $n$-th permutation, $\tilde{\Delta}^{(n)}=\lvert\sum_{i=1}^{k}\phi(\tilde{x}_{i}^{(n)})/k-\sum_{i=1}^{k}\phi(\tilde{y}_{i}^{(n)})/k\rvert$;
@@
In the experiment, we consider the composite hypothesis testing, which means the null hypothesis is $H_{0}:\lvert\phi(x_{t})-\phi(y_{t})\leq\rvert\epsilon$. If we still use the above permutation test, it will become much easier for $\Delta\geq\tilde{\Delta}^{(n)}$ to hold, even when $H_{0}$ is actually true. This would result in significantly higher FPRs. Thus, we only check p-values when the observed $\Delta$ exceeds the estimated $\epsilon$. The rejection time and FPR that we plot are the average values across 1000 repeated runs for each significance level.
@@
As observed in Figure 7, the permutation tests without correction always have higher FPRs under $H_{0}$ than that with corrected significance levels. This phenomenon aligns with the fact that without significance level adjustments, it is impossible to control the type-I error. Permutation tests with large batch sizes demonstrate relatively low FPRs, which are approximate equal to 0. However, this seemingly excellent performance is due to the fact that the preset $\epsilon$ values are much larger than the actual absolute difference in mean scores between sequences of XSum texts and 2024 Olympic news. This discrepancy results in fewer or no p-value checks, thus sustaining $H_{0}$. Consequently, the FPRs are nearly identical across each significance level $\alpha$. Permutation tests are sensitive to the discrepancies between two sequences, and the $\Delta$ value tends to change more with smaller batch sizes due to variation among scores of texts from the same source. Although a permutation test can reject $H_{0}$ after the first batch test, it consistently exhibit an FPR greater than $\alpha$. Moreover, even if we set the value of $\epsilon$ based on a much larger sample size, rather than from estimates derived from a few points, there will still be variance between the $\Delta$ value calculated in batches and the preset $\epsilon$. As long as $\Delta$ calculated by a batch of samples is greater than the preset $\epsilon$, the permutation test is likely to have a large FPR under $H_{0}$. Compared to fixed-time methods, our method can use parameter estimates based on just a few points to ensure faster rejection and lower FPRs. It can save time with high acuuracy especially when there is no prior knowledge of the threshold $\epsilon$ for composite hypotheses.
@@
(a) Comparisons between our method and the permutation test without correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(b) Comparisons between our method and the permutation test with correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(c) Comparisons between our method and the permutation test without correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(d) Comparisons between our method and the permutation test with correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(e) Comparisons between our method and the permutation test without correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2 with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(f) Comparisons between our method and the permutation test with correction for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2 with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.

### v3

@@ Appendix G Experimental Results of Detecting 2024 Olympic News or Machine-generated News
Experimental Results for Two Scenarios. Figure 5 and Figure 6 show the results of detecting real Olympic news and news generated by Gemini-1.5-Flash, Gemini-1.5-Pro, and PaLM 2, when designating human-written text from the XSum dataset as $x_{t}$ in Scenario 1 and Scenario 2, respectively. In Scenario 1, our algorithm consistently controls the FPRs below the significance level $\alpha$ for all source models, scoring models, and score functions. This is because we set the true value of $\Delta$ between two sequences of human texts to $\epsilon$, which satisfies the null hypothesis condition, i.e., $|\mu_{x}-\mu_{y}|\leq\epsilon$. This ensures that the wealth remains a supermartingale. Texts generated by PaLM 2 are detected almost immediately by most score functions within 100 time steps, as illustrated in Figure 5(e) and Figure 5(f). Conversely, fake Olympic news generated by Gemini-1.5-Pro often fails to be identified as LLM-generated within 500 time steps under certain score functions, as shown in Figure 5(c) and Figure 5(d). Vertical lines in Figure 5(a)-5(d) are displayed because the the $\Delta$ values for human texts and fake news, as shown in Table 1, are smaller than the corresponding $\epsilon$ values. This means that the score discrepancies between fake news and XSum texts do not exceed the threshold necessary for rejecting $H_{0}$. Although under $H_{1}$, the $\Delta$ for Entropy when using Gemma-2B to score texts generated by Gemini-1.5-Pro is 0.2745, larger than the value of $\epsilon$ ($0.2690$), the discrepancy is too small to lead to a rejection of $H_{0}$ within 500 time steps.
@@
(a) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Flash), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(b) Results for detecting the source $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Flash), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(c) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Pro), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(d) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Pro), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(e) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by PaLM 2), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(f) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by PaLM 2), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
Figure 5: Results for detecting 2024 Olympic news and machine-generated news with Algorithm 3 for Scenario 1. We use 3 source models: Gemini-1.5-Flash, Gemini-1.5-Pro and PaLM 2 to generate fake news and 2 scoring models: Neo-2.7, Gemma-2B. The left column displays results using the Neo-2.7 scoring model, while the right column presents results using the Gemma-2B scoring model. Score functions of supervised classifiers (RoB-base and RoB-large) are independent of scoring models.
(a) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Flash), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(b) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Flash), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(c) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Pro), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(d) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by Gemini-1.5-Pro), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
(e) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by PaLM 2), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Neo-2.7.
(f) Results for detecting the source of $y_{t}$ (2024 Olympic news vs. news generated by PaLM 2), with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
Table 1: Values of $\Delta$, which are calculated according to $\Delta=\left\lvert{\sum_{i=1}^{500}\phi(x_{i})}/{500}-{\sum_{j=1}^{500}\phi(y_{j})}/{500}\right\rvert$, where $\phi(x_{i})$ is score of the $i$-th text $x_{i}$ from XSum, $\phi(y_{i})$ is score of the $j$-th text $y_{j}$ from the source to be detected. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ for each test scenario. For instance, in calculating $\Delta$ for the third column, $y_{j}$ represents the $j$-th fake news generated by Gemini-1.5-Flash based on pre-tokens of Olympic 2024 news. For the fourth column, $y_{j}$ refers to the $j$-th 2024 Olympic news articles. Values in Column “Human, Human" are also used to set $\epsilon$ values for tests in Scenario 1.
@@
Table 2: Values of $d_{t}$ used in Scenario 1, for which we assume that the range of $|g_{t}|=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ is known beforehand. Specifically, $d_{t}$ is calculated as $\max_{i,j\leq 500}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by Gemini-1.5-Flash, prompted by pre-tokens of 2024 Olympic news. This calculation ensures that $d_{t}\geq\lvert\phi(x_{t})-\phi(y_{t})\rvert$ for any time point $1\leq t\leq 500$. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column Similarly, the fourth column calculates the maximum difference between the scores of all XSum texts and texts sample of 2024 Olympic news.. The derived $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},0]$.
@@
Moreover, we found that the rejection time is related to the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$. According to the definition of nonnegative wealth $W_{t}^{A}=W_{t-1}^{A}(1-\theta_{t}(g_{t}-\epsilon))$ or $W_{t}^{B}=W_{t-1}^{B}(1-\theta_{t}(-g_{t}-\epsilon))$, large $-\theta_{t}$ within the range $[0,1/2d_{t}]$ will result in large wealth which allows to quickly reach the threshold for wealth to correctly declare the unknown source as an LLM. Based on the previous proposition of the expected time upper bound for composite hypothesis, we guess that the actual rejection time in our experiment is probably related to the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$, where $d_{t}$ is a certain value for any $t$ in each test as shown in Table 2. We define the relative magnitude as $\frac{\Delta-\epsilon}{d_{t}-\epsilon}$ and sort the score functions by this ratio from largest to smallest for Scenario 1, as displayed in the Rank column in Table 3. This ranking roughly corresponds to the chronological order of rejection shown in Figure 5. The quick declaration of an LLM source when $y_{t}$ is generated by PaLM 2 and the slower rejection of $H_{0}$ when $y_{t}$ is generated by Gemini-1.5-Pro in Figure 5 could be attributed to the relatively larger and smaller values of $\frac{\Delta-\epsilon}{d_{t}-\epsilon}$, respectively. The negative ratios arise because $\Delta<\epsilon$, leading to the vertical lines in the figures. For Scenario 2, the values of $\frac{\Delta-\epsilon}{d_{t}-\epsilon}$ are provided in Table 4, and the conclusions from Scenario 1 extend to Scenario 2.
@@
Table 3: Values of the ratio $\frac{\Delta-\epsilon}{d_{t}-\epsilon}$ for Scenario 1, where $\Delta$ and $\epsilon$ are listed in Table 1, $d_{t}$ are shown in Table 2. We sort the score function according to the ratio from largest to smallest, as shown in the Rank column. This ranking roughly corresponds to the chronological order of rejection in in Figure 5.
@@
Table 4: Values of the ratio $\frac{\Delta-\epsilon}{d_{t}-\epsilon}$ for Scenario 2, where $\Delta$ and $\epsilon$ are listed in Table 1 and Table 5 respectively, $d_{t}$ are shown in Table 6. We sort the scoring function according to the ratio from largest to smallest, as shown in the Rank column. This ranking roughly corresponds to the chronological order of rejection in in Figure 6.
@@
Table 5: Average values of $\epsilon$ used in Scenario 2 estimated by 20 texts in sequence of human-written text $x_{t}$. Every two columns starting from the third column represent the $\epsilon$ values used for $H_{1}$ and $H_{0}$ for each test scenario. For instance, the third column calculates $\epsilon$ for tests between XSum text and Gemini-1.5-Flash-generated text sequences by scoring 20 XSum texts, dividing them into two equal groups, and then doubling the average absolute mean difference between these groups across 1000 random shuffles. The fourth column follows the same method to determine the $\epsilon$ value for tests between XSum texts and 2024 Olympic news. This is calculated as $\epsilon=2\cdot\frac{1}{1000}\sum_{n=1}^{1000}\left\lvert{\sum_{i=1}^{10}\phi(x_{i}^{(n)}})/{10}-{\sum_{i=11}^{20}\phi(x_{i}^{(n)}})/{10}\right\rvert$, where $\phi(x_{i}^{(n)})$ denotes the score of the $i$-th text after the $n$-th random shuffling of 20 text scores.
@@
Table 6: Average values of $d_{t}$ used in Scenario 2 estimated by using the first 10 texts of each sequence. Every two columns starting from the third column represent the $d_{t}$ values used for $H_{1}$ and $H_{0}$ for each test scenario. Specifically, for the third column, we get the first 10 samples from XSum and the first 10 observed texts generated by Gemini-1.5-Flash. We then calculate the maximum difference between $\phi(x_{i})$ and $\phi(y_{j})$ for any $1\leq i\leq 10$, $1\leq j\leq 10$. We double this maximum value to estimate $d_{t}$ value, i.e., $d_{t}=2\cdot\max_{i,j\leq 10}\lvert\phi(x_{i})-\phi(y_{j})\rvert$. The fourth column follows a similar calculation for detecting 2024 Olympic news with XSum texts, where $\phi(x_{i})$ represents score of $i$-th text from XSum, $\phi(y_{j})$ denotes the score of the $j$-th text of 2024 Olympic news.
@@
Based on the results, when we know the actual value of $d_{t}$ and $\epsilon$, our algorithm is very effective. When we estimate their values based on the previous samples that we get, the algorithm can still exhibit a good performance for most score functions. It can be inferred that the rejection time and FPRs are effected by the score function $\phi(\cdot)$ and the scoring model that we select. If the configuration can further amplify the score discrepancy between human-written texts and machine-generated texts, the rejection time can be shortened. If the the score discrepancy between human texts is small, the value FPR will be low.
Comparisons with Baselines. Permutation test is a fixed-time test. The goal is to test whether the source of text $y_{t}$ is the same as that of the human-written text $x_{t}$, i.e., whether their scores are from the same distribution. If we choose the mean value as the test statistic, the null hypothesis is that the means are equal ($H_{0}:\mu_{x}=\mu_{y}$) with a batch size of $k$. Once we have generated $k$ samples from these two sources, we conduct the test. When $H_{0}$ is true, the samples are drawn from the same distribution, which means that the observed discrepancy between the two batches of scores is supposed to be minimal. The test determines whether this difference between the sample means is large enough to reject $H_{0}$ at a significance level. Specifically, we compare the p-value with the significance level $\alpha$ for each batch in an uncorrected test, and $\alpha/2^{j}$ for the $j$-th batch in a corrected test. The permutation test is conducted as below:
(1) Calculate the observed mean difference $\Delta=\left\lvert\sum_{i=1}^{k}\phi(x_{i})/k-\sum_{j=1}^{k}\phi(y_{j})/k\right\rvert$ of these two batches, and assume that $H_{0}$ is true;
(2) Combine these two sequences into one dataset, reshuffle the data, and divide it into two new groups. This is the permutation operation. Calculate the sampled absolute mean difference for the $n$-th permutation, $\tilde{\Delta}^{(n)}=\lvert\sum_{i=1}^{k}\phi(\tilde{x}_{i}^{(n)})/k-\sum_{j=1}^{k}\phi(\tilde{y}_{j}^{(n)})/k\rvert$;
@@
In the experiment, we consider the composite hypothesis testing, which means the null hypothesis is $H_{0}:\lvert\mu_{x}-\mu_{y}\rvert\leq\epsilon$. If we still use the above permutation test, it will become much easier for $\Delta\geq\tilde{\Delta}^{(n)}$ to hold, even when $H_{0}$ is actually true. This would result in significantly higher FPRs. Thus, we only check p-values when the observed $\Delta$ exceeds the estimated $\epsilon$. The rejection time and FPRs that we plot are the average values across 1000 repeated runs under each significance level $\alpha$.
@@
As observed in Figure 7, the permutation tests without correction always have higher FPRs under $H_{0}$ than that with corrected significance levels. This phenomenon aligns with the fact that without significance level adjustments, it is impossible to control the type-I error. Permutation tests with large batch sizes demonstrate relatively low FPRs, which are approximately equal to 0. However, this seemingly excellent performance is due to the fact that the preset $\epsilon$ values are much larger than the actual absolute difference in mean scores between sequences of XSum texts and 2024 Olympic news. This discrepancy results in fewer or no p-value checks, thus sustaining $H_{0}$. Consequently, the FPRs are nearly identical across each significance level $\alpha$. Permutation tests are sensitive to the discrepancies between two sequences, and the $\Delta$ value tends to change more with smaller batch sizes due to variation among scores of texts from the same source. Although a permutation test can reject $H_{0}$ after the first batch test, it consistently exhibit an FPR greater than $\alpha$. Moreover, even if we set the value of $\epsilon$ based on a much larger sample size, rather than from estimates derived from a few points, there will still be variance between the $\Delta$ value calculated in batches and the preset $\epsilon$. As long as $\Delta$ calculated by a batch of samples is greater than the preset $\epsilon$, the permutation test is likely to have a large FPR under $H_{0}$. Compared to fixed-time methods, our method can use parameter estimates based on just a few points to ensure faster rejection and lower FPRs. It can save time with high acuuracy especially when there is no prior knowledge of the threshold $\epsilon$ for composite hypotheses.
@@
(a) Comparisons between our method and the permutation test without correction, $y_{t}$ are 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(b) Comparisons between our method and the permutation test with correction, $y_{t}$ are 2024 Olympic news or news generated by Gemini-1.5-Flash, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(c) Comparisons between our method and the permutation test without correction, $y_{t}$ are 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(d) Comparisons between our method and the permutation test with correction, $y_{t}$ are 2024 Olympic news or news generated by Gemini-1.5-Pro, with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(e) Comparisons between our method and the permutation test without correction, $y_{t}$ are 2024 Olympic news or news generated by PaLM 2 with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.
@@
(f) Comparisons between our method and the permutation test with correction, $y_{t}$ are 2024 Olympic news or news generated by PaLM 2 with human-written text $x_{t}$ sampled from XSum. The scoring function used is Neo-2.7.